A police station had to deploy a police officer for an emergency multiple times in the last four evenings. The table below shows the number of emergencies each evening. (Round your answers to 3 decimal places.) Forecast for tomorrow (calls) Weekday Number of calls each day 10 Monday Tuesday Wednesday Thursday 7 2 7 Complete the Forecast for tomorrow column in the table above. Use an exponential smoothing a. forecasting approach, with a-0.2 and a forecast for Monday of 10. (Round your answer to 3 decimal places.) smoothing forecasting approach? b. What would be their forecast for Friday using the same exponential
Answer #1

A) Exponential smoothing method with Alpha = 0.2: The formula to compute the forecast using the exponential smoothing method is

Current forecast = Previous Forecast + [Alpha x (Previous Actual – Previous Forecast)

Given that Forecast for Monday is 10.

Forecast_Tue = Forecast_Mon + [Alpha x (Actual_Mon – Forecast_Mon)]

= 10 + [0.20 x (7 – 10)] = 9.400

Forecast_Wed = Forecast_ Tue + [Alpha x (Actual_ Tue – Forecast_ Tue)]

= 9.400 + [0.20 x (2.000 – 9.400)] = 7.920

Forecast_Thu = Forecast_ Wed + [Alpha x (Actual_ Wed – Forecast_ Wed)]

= 7.920 + [0.20 x (7 – 7.920)] = 7.736

B) Forecast_Fri = Forecast_ Thu + [Alpha x (Actual_ Thu – Forecast_ Thu)]

= 7.736 + [0.20 x (11 – 7.736)] = 8.389